OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -51 A.-126 B.-123 C.-120 D.-117

OpenStudy (anonymous):

\[T _{n}=a+(n-1)d\] \[T_{1}=13 \[T_{17}=13+(17-1)d\] \[-51=13+16d\] \[16d=-51-13\] \[16d=-64\] \[d=-4\] D is the common difference. Now you have the common difference and you were already given the first term, you can now find the 35th term using the formula for finding arithmetic terms.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Do you know the formula: \[T_{n}=a+(n-1)d\]

OpenStudy (anonymous):

Do you know this formula?

OpenStudy (anonymous):

no .

OpenStudy (anonymous):

Then how are you supposed to answer question without knowing the formulas first?

OpenStudy (anonymous):

answer the*

OpenStudy (anonymous):

i really dont know , math is not my thing .

OpenStudy (anonymous):

Then sorry mate, but you need to pay attention in class. If you do that then you wouldn't be saying math is not my thing. If you do pay attention but you don't understand the concepts in class, why wouldn't you raise your hand or go to the teacher and ask for help?

OpenStudy (anonymous):

im not that smart , i have challenges..

OpenStudy (anonymous):

is the answer b ?

OpenStudy (anonymous):

Why are you saying you're not smart and that you have challenges while you just go on and give me the correct answer? Shisssh....